Find four consecutive even integers such that the sum of the second and fourth be 74.
Answer by RAY100(1637)

Let consecutive integers be,,,x,,,x+2,,,x+4,,,,x+6
.
then,,,(x+2) +(x+6) =74,,,,,,,,2x+8=74,,,,,x=33
.
Integers are, 33,35,37,39,,,,and sum of 2nd & 4th is 35+39=74
.
However these are odd,, not even integers.
.